# online_appendix

### v2

@@ Abstract
Keywords: international trade, low-value transactions, fast fashion, e-commerce, online retail platforms, Spain.
1 Introduction
The new stage of the globalization process that began in the 1980s presents some characteristics, such as the revolution in information and communication technology, the emergence of China as a global manufacturing exporter, and the expansion of global value chains, which were unknown in previous integration episodes (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). This paper documents an additional novel feature of this globalization stage: the increase in the share of low-value transactions in the total number of international trade transactions.
We define a low-value transaction as the exchange of a product between a domestic and a foreign trade operator (i.e., a firm or an individual) whose value does not exceed 150 euros. This value is the de minimis threshold used by the European Union (EU): any shipment from a non-EU country to a EU country whose value is equal to or less than 150 euros does not pay any tariffs when entering the EU. [^1] Article 23 of the Council Regulation 1186/2009. The Council Regulation No. 918/83 that established the de minimis threshold in the EU denoted transactions equal to or below the de minimis threshold as consignments of negligible value. Using the universe of Spanish international trade transactions, this paper shows that the share of low-value transactions in the total number of transactions increased from 9% to 61% in exports and from 14% to 54% in imports between 1997 and 2023.
@@ 1 Introduction
We offer two explanations for the increase in the share of low-value transactions in the total number of international trade transactions. First, the increase in the number of low-value transactions in exports until 2021 is explained by clothing and footwear firms. During the last decades, these firms have developed a so-called fast fashion strategy, characterized by the rapid adaptation of clothing displayed in the store to changing consumer preferences and the creation of scarcity with collections that end quickly. This strategy requires a frequent restocking of stores, which manifests itself in a large number of low-value export transactions between the distribution center in the home country and the stores located abroad. This frequency is higher in destinations geographically far from the distribution center, as products need more time to reach the destination. In addition, products shipped to distant destinations are more likely to be transported by air to ensure a rapid response to a change in preferences. Consistent with this argument, we show that the leading fashion retailers explained more than 60% of the increase in the number of low-value export transactions between 1997 and 2021. Furthermore, we find that the largest number of low-value export transactions were destined to countries geographically far from Spain and products were transported by air.
@@
Second, the increase in the number of low-value import transactions is related to the emergence of online retail platforms, such as Shein or Temu, that allow consumers to purchase a wide range of products while providing foreign manufacturers with access to a global market. Transactions mediated through these platforms are denoted as direct to consumer, because purchases go directly from the overseas producer to the local consumer (Fajgelbaum and Khandelwal,, 2024). Consistent with this explanation, we show that the increase in the number of low-value import transactions rocketed between 2021 and 2023, when the above-mentioned platforms became popular. Furthermore, these import transactions originated in China, where these online platforms were founded. The geographical distribution and administrative justification for the large number of low-value import transactions from China are also consistent with a direct-to-consumer shipment explanation. For earlier periods, we provide evidence consistent with the argument that online stores and marketplaces explain the increase in the number of low-value import transactions. Finally, we show that the increase in the number of low-value export transactions occurring from 2021 onward, which is not related to fast-fashion retailers, is also consistent with the emergence and development of online marketplaces.
@@
Is the increase in the share of low-value transactions in the total number of trade transactions a particular feature of Spanish trade, or is it a global trend? Additional analyses seem to confirm the trend observed for Spain, at least for imports. Fajgelbaum and Khandelwal, (2024) show that the volume of de minimis imports in the United States multiplied by 10 between 2012 and 2013, and the share of de minimis imports in the total value of US imports increased from 0.002% to 1.720% during the same period. In the case of the EU, according to European Commission, (2023), the volume of e-commerce imports from countries outside the EU tripled between 2019 and 2022. Regarding exports, the fact that some Spanish firms are global leaders in the fashion industry could have enhanced the share of low-value transactions in the volume of export transactions.
@@
The increase in the share of low-value transactions in the total number of transactions has important implications for international trade. First, a growing share of imports enter the US and the EU with a value equal to or below the de minimis threshold to avoid paying tariffs is argued. This concern is particularly strong in the US since tariffs on many Chinese goods were increased in 2018. Furthermore, there are claims that de minimis enables the import of goods that do not meet health and safety standards. [^2] See “Where Textile Mills Thrived, Remnants Battle for Survival” by Jordyn Holman, The New York Times, 21 January 2024 at https://www.nytimes.com/2024/01/21/business/economy/textile-mills-carolina-trade-de-minimis.html and “The EU targets China’s Temu and Shein with proposed import duty” by Andy Bounds Paola Tamma, Financial Times, 3 July 2024 at https://www.ft.com/content/1c4c0bee-f67e-404b-877d-e0cb38faf2d6. The alleged misuse of the de minimis threshold has led to proposals to eliminate it both in the United States and in the EU. Therefore, the increase in the share of low-value transactions could have important implications for trade policy.
@@
Second, the increase in the share of low-value transactions has been facilitated by the emergence of e-commerce platforms that allow consumers to choose from a wide range of products. These platforms allow consumers to be directly involved in international trade, and, therefore, their purchasing decisions have a greater weight in shaping trade flows (Volpe-Martincus et al.,, 2024). If there are differences in the variables that govern consumer purchasing decisions from those of firms, it will be important to take those differences into account when designing economic policies. For example, Volpe-Martincus et al., (2024) show that low-value transactions are not affected by trade agreements and this paper shows that the time-sensitiveness of low-value transactions leads to a greater use of air transport.
@@
Third, low-income consumers import more low-value products than high-income ones (Fajgelbaum and Khandelwal,, 2024). The increase in low-value import transactions leads to a greater direct participation of low-income consumers in trade and underscores the need to consider the impact of trade policies, such as the elimination of the de minimis threshold, on income inequality.
@@
To the best of our knowledge, this paper is the first to document the increase in the share of low-value transactions in the total number of international trade transactions. This discovery adds to the literature that has analyzed the features of the globalization stage that began in the 1980s (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). We also add to the literature on the role of the de minimis threshold in international trade. As mentioned above, Fajgelbaum and Khandelwal, (2024) document the rise in de minimis imports in the US and show that less affluent consumers are more likely to be recipients of low-value imports. We contribute to this literature by analyzing both low-value export and import transactions, quantifying the share in value and volume of low-value transactions in total trade transactions, exploring a longer time period, and highlighting some stylized facts about low-value transactions. We also offer a new explanation for the increase in the number of low-value transactions: the fast-fashion strategy followed by fashion retailers.
@@
Finally, we also contribute to the literature exploring the relationship between fast fashion and international trade. Fernandes and Tang, (2020) show that Portuguese clothing manufacturers responded to Chinese competition by increasing the frequency of their shipments. We show that the high frequency of shipments by fashion retailers played a very important role in the increase in the number of low-value export transactions. Furthermore, we show that these transactions were concentrated in countries that were geographically far from Spain and products that were transported by air. These latter findings also link our paper to the literature that analyzes how the time sensitivity of demand determines the choice of transport means (Hummels and Schaur,, 2013) and the lumpiness of trade (Alessandria et al.,, 2010; Kropf and Sauré,, 2014; Hornok and Koren, 2015a, ; Hornok and Koren, 2015b, ; Békés et al.,, 2017).
@@ 2 Data and stylized facts
We obtain trade data on goods for the period 1997-2023 from the Customs and Excise Department of the Spanish Tax Agency (AEAT-Customs). Each observation captures the exchange of a product between a Spanish trade operator (a firm or an individual) and a foreign trade operator. It provides information on the direction of the trade flow (exports or imports), the Spanish trade operator’s customs identification code, the product in the 8-digit Combined Nomenclature (CN) classification, the value (in euros), the country of destination or origin, the Spanish province of destination or origin, the month in which the transaction occurred, and the date the transaction was documented in customs. It does not provide an identification of the foreign partner. We remove all transactions that do not involve the transfer of property in exchange for financial compensation. Hence, our dataset does not include transactions related to gifts, donations, or return of goods. [^3] We exclude Andorra and Gibraltar from the list of origins and destinations because, due to their location, trade with these territories is similar to domestic sales for Spanish firms.
@@
The AEAT-Customs database combines information from two sources: Intrastat and Extrastat. Intrastat captures firms’ trade flows with EU partners. In 2023, firms had to present a monthly Intrastat declaration if they accumulated an export (import) value with EU members of 400,000 euros in the current year or in the previous year. Extrastat captures firms’ trade flows with non-EU countries. This information is based on the customs declaration EU-based firms must present every time they perform an export or import operation with a non-EU country. Firms must present a customs declaration regardless of the amount of trade they accumulated in the previous year or have accumulated in the current year. Due to Intrastat’s data presentation threshold, AEAT-Customs does not capture all the small- and large-value transactions occurring between Spanish firms and their EU partners. This should not affect our estimations, because the below-the-Intrastat threshold trade only represented 1.8% of total trade. [^4] This data, estimated by AEAT-Customs, corresponds to 2022 (the last year with definitive data). It is available at: https://sede.agenciatributaria.gob.es/Sede/estadisticas/estadisticas-comercio-exterior/principales-resultados-serie/2023.html. In any case, the appendix shows that there are no major differences in the evolution of the number of low-value trade transactions between Spain and EU countries and Spain and non-EU countries.
@@ C. Top 20 HS 2-digit chapters by share in total number of low-value transactions (2023)
The first stylized fact, whose documentation constitutes the main contribution of this paper, is the large increase in the share of low-value transactions in the total number of international trade transactions. It rises from 9% to 61% in exports and from 14% to 54% in imports. That is, in 2023, low-value transactions represented more than half of the total number of trade transactions. [^5] The median values of an export and import transaction in 1997 were 4,678 and 2,776 euros, respectively. In 2023, they decreased to 77 and 114 euros. As shown in Figure A.1 in the Appendix, this stylized fact is robust to using a low-value threshold adjusted for inflation, a low-value threshold based on the 800 USD de minimis value applied by the United States, and to splitting the sample into EU and non-EU members. [^6] The exception is that the share of low-value export transactions in the total number of transactions between 2021 and 2023 decreased in non-EU27 countries and increased in EU27 countries.
As mentioned above, in general, a transaction cannot be equated with a shipment. However, if some conditions are met, a transaction can be classified as a shipment. [^7] This happens when for a given export (or import) transaction, we cannot find another transaction that is performed by the same firm, happens in the same month, uses the same mode of transport, originates in the same province, is destined to the same country, and was recorded in customs on the same date. Although the percentage of transactions that can be classified as shipments is small, the evolution of the share of low-value shipments in total shipments between 1997 and 2023 is qualitatively similar to that found for transactions. [^8] The share of low-value shipments in total shipments multiplies by 3 both for exports and imports between 1997 and 2023.
@@
The total value of low-value exports increased from 16 million euros in 1997 to 1,478 million euros in 2023 (multiplied by 92). For imports, the value of low-value transactions increased from 30 million euros to 735 million euros (multiplied by 24). These increases were much larger than those in total value, which multiplied by 4 for both exports and imports between 1997 and 2023. Despite these large relative increases, our second stylized fact is that low-value transactions still represent a very small share of the total value of trade: 0.4% and 0.2% of the total value of exports and imports in 2023, respectively (panel B). [^9] In 1997, the figures were 0.02% and 0.03% for exports and imports, respectively. Fajgelbaum and Khandelwal, (2024) calculated that de minimis imports represented 7.2% of US imports of consumer goods in 2023. If we used the 800 USD threshold to define low-value transactions, these would represent 2.6% of the Spanish imports of consumer goods in 2023. For exports, the share of low-value transactions would increase to 3.8%. Therefore, when using the US de minimis threshold, low-value transactions represent a nonnegligible share of the Spanish trade in consumer goods. [^10] Using the US thresholds, low-value transactions would represent 1.5% and 0.8% of the total value of exports and imports in 2023.
@@
A major limitation of our dataset is that most low-value transactions do not report the traded product. Specifically, 70% and 71% of low-value trade transactions in 2023 did not report the exported or imported product, respectively. This occurs because export transactions equal to or below the de minimis value of the destination country, and import transactions equal to or below 150 euros are not obliged to report the traded product. Keeping this limitation in mind, the third stylized fact is that clothing and footwear represented the largest share of low-value transactions that recorded the traded product in 2023: 41% and 30% of the total number of low-value export and import transactions, respectively (panel C of Figure 1). By broad economic activities, consumption goods represented 67% and 53% of the total number of low-value export and import transactions, respectively, in 2023; intermediate goods represented 26% and 37% of export and import transactions, respectively; and parts related to capital goods represented 7% and 10% of export and import transactions, respectively. Interestingly, in 1997, books were the top product in exports and electrical machinery in imports (panel A of Figure A.2 in the appendix). The share of clothing and footwear in 1997 was 12% and 13% for exports and imports, respectively. This comparison highlights that the clothing and footwear industries played an important role in the growth of the number of low-value transactions. [^11] Panel A of Figure A.3 in the appendix shows that clothing and footwear also occupied prominent roles among the most exported and imported products in high-value transactions in 2023. However, their share in the total number of transactions was less than that in low-value transactions: 23% in large-value export transactions versus 41% in low-value transactions, and 22% versus 30% in import transactions.

### v3

@@ Abstract
Keywords: international trade, low-value transactions, de minimis, e-commerce, fast fashion, online retail platforms, Spain.
The new stage of the globalization process that began in the 1980s presents some characteristics, such as the revolution in information and communication technologies, the emergence of China as a global manufacturing exporter, and the expansion of global value chains, which were unknown in previous integration episodes (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). This paper documents an additional novel feature of this globalization stage: the increase in the share of low-value transactions in the total number of international trade transactions.
We define a low-value transaction as the exchange of a product between a domestic and a foreign trade operator (i.e., a firm or an individual) whose value does not exceed 150 euros. This value is the de minimis threshold applied in the European Union (EU): any shipment dispatched to its recipient containing goods of a total value not exceeding 150 euros should be admitted free of import duties. [^1] Article 27 of the Council Regulation No. 918/83 and Article 1.3 of the Council Regulation 1186/2009. Using the universe of Spanish international trade transactions, this paper shows that the share of low-value transactions in the total number of transactions increased from 9% to 61% in exports and from 14% to 54% in imports between 1997 and 2023.
@@
We offer two explanations for the increase in the number of low-value transactions. First, the increase in the number of low-value transactions in exports until 2021 is explained by clothing and footwear firms. During the last decades, these firms have developed a so-called fast fashion strategy, characterized by the rapid adaptation of clothing displayed in the store to changing consumer preferences and the creation of scarcity with collections that end quickly. This strategy requires a frequent restocking of stores, which manifests itself in a large number of low-value export transactions between the distribution center in the home country and the stores located abroad. This frequency is higher in destinations geographically far from the distribution center, as products need more time to reach the destination. In addition, products shipped to distant destinations are more likely to be transported by air to ensure a rapid response to a change in preferences. Consistent with this argument, we show that the leading fashion retailers explained more than 60% of the increase in the number of low-value export transactions between 1997 and 2021. Furthermore, we find that the largest number of low-value export transactions were destined to countries geographically far from Spain and products were transported by air.
@@
Second, the increase in the number of low-value import transactions is related to the popularization of online retail platforms, such as Shein or Temu, that allow consumers to purchase a wide range of products while providing foreign manufacturers with access to a global market. Transactions mediated through these platforms are denoted as direct to consumer, because purchases go directly from the overseas producer to the local consumer. Consistent with this explanation, we show that the increase in the number of low-value import transactions rocketed between 2021 and 2023, when the above-mentioned platforms became popular. Furthermore, these import transactions originated in China, where these online platforms were founded. The geographical distribution and administrative justification for the large number of low-value import transactions from China are also consistent with a direct-to-consumer shipment explanation. For earlier periods, we provide evidence consistent with the argument that online stores and marketplaces explain the increase in the number of low-value import transactions. Finally, we show that the increase in the number of low-value export transactions occurring from 2021 onward, which is not related to fast-fashion retailers, is also consistent with the emergence and development of online marketplaces.
@@
Is the increase in the share of low-value transactions in the total number of trade transactions a particular feature of Spanish trade, or is it a global trend? Other studies seem to confirm the trend observed for Spain, at least for imports. Fajgelbaum and Khandelwal, (2024) show that the volume of de minimis imports in the United States multiplied by 10 between 2012 and 2023, and the share of de minimis imports in the total value of US imports increased from 0.002% to 1.720% during the same period. In the case of the EU, according to European Commission, (2023), the volume of E-commerce imports from countries outside the EU tripled between 2019 and 2022. Regarding exports, the fact that some Spanish firms are global leaders in the fashion industry could have enhanced the share of low-value transactions in the volume of export transactions, and it is reasonable to expect the share to be lower in other countries.
@@
The increase in the share of low-value transactions in the total number of transactions has important implications for international trade. First, a growing share of imports enter the US and the EU with a value equal to or below the de minimis threshold to avoid paying tariffs is argued. This concern is particularly strong in the US since tariffs on many Chinese goods increased in 2018.  [^2] See “Where Textile Mills Thrived, Remnants Battle for Survival” by Jordyn Holman, The New York Times, 21 January 2024 at https://www.nytimes.com/2024/01/21/business/economy/textile-mills-carolina-trade-de-minimis.html and “The EU targets China’s Temu and Shein with proposed import duty” by Andy Bounds Paola Tamma, Financial Times, 3 July 2024 at https://www.ft.com/content/1c4c0bee-f67e-404b-877d-e0cb38faf2d6. In response to these concerns, the US Government announced on September 13, 2024, that it would eliminate the de minimis exception on approximately 40% of US imports and the European Commission opened a formal proceeding against Temu on October 31, 2024.333The US Government statement is available at: https://www.whitehouse.gov/briefing-room/statements-releases/2024/09/13/fact-sheet-biden-harris-administration-announces-new-actions-to-protect-american-consumers-workers-and-businesses-by-cracking-down-on-de-minimis-shipments-with-unsafe-unfairly-traded-products/. The European Commission statemet is available at:https://ec.europa.eu/commission/presscorner/detail/en/ip_24_5622 These decisions highlight how the increase in the volume of low-value transactions could trigger important changes in trade policy. [^4] For example, as explained by Fajgelbaum and Khandelwal, (2024), between June-September 2024, Brazil, Chile, Philippines, and Turkey have either reduced the de minimis threshold, imposed tariffs on low-value transactions, or introduced new customs fees for the processing of small transactions.
Second, the increase in the share of low-value transactions has been facilitated by the emergence of E-commerce platforms that allow consumers to choose from a wide range of products. This fact underscores the crucial role that the high-speed Internet and mobile applications play in enabling and promoting international trade. However, mobile apps may also gain access to extensive user data and governments are concerned that this puts individuals’ privacy and national security at risk. Consequently, low-value transactions may emerge as a further point of contention in the ongoing trade versus security debate.
@@
Third, E-commerce platforms allow consumers to be directly involved in international trade, and therefore their purchasing decisions have a greater weight in shaping trade flows (Volpe-Martincus et al.,, 2024). If there are differences in the variables that govern consumer purchasing decisions from those of firms, it will be important to take those differences into account when designing economic policies. For example, Volpe-Martincus et al., (2024) show that imports by individuals are not affected by trade agreements, and this paper reveals that the time sensitiveness of low-value transactions leads to a greater use of air transport.
@@
Fourth, low-income consumers import more low-value products than high-income ones (Fajgelbaum and Khandelwal,, 2024). This underscores the need to consider the impact of trade policies, such as the elimination of the de minimis threshold, on income inequality.
Fifth, many low-value goods are shipped directly from manufacturers to consumers. These shipments are transported by air and save warehouse costs, introducing a major change in the logistics of international trade. Finally, the use of air transport suggests that low-value transactions may have a greater environmental impact than the average firm-to-firm transaction (Biancolin and Rotaris,, 2024). In addition, the environmental impact of low-value transactions could be aggravated by the fact that many E-commerce purchases are returned (Zhang et al.,, 2022).
@@
To the best of our knowledge, this paper is the first to document the increase in the share of low-value transactions in the total number of international trade transactions. This discovery adds to the literature that has analyzed the features of the globalization stage that began in the 1980s (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). We also add to the literature on the role of the de minimis threshold in international trade. Fajgelbaum and Khandelwal, (2024) document the increase in de minimis imports in the US and show that less affluent consumers are more likely to be recipients of low-value imports. We contribute to this literature by analyzing both low-value export and import transactions, quantifying the share in value and volume of low-value transactions in total trade transactions, exploring a longer time period, and highlighting some stylized facts about low-value transactions. We also offer a new explanation for the increase in the number of low-value transactions: the fast-fashion strategy followed by fashion retailers.
@@
We also contribute to the literature exploring the relationship between fast fashion and international trade. Fernandes and Tang, (2020) show that Portuguese clothing manufacturers responded to Chinese competition by increasing the frequency of their shipments. We show that the high frequency of shipments by fashion retailers played a very important role in the increase in the number of low-value export transactions. Furthermore, we show that these transactions were concentrated in countries that were geographically far from Spain and products that were transported by air. These latter findings also link our paper to the literature that analyzes how the time sensitivity of demand determines the choice of transport means (Hummels and Schaur,, 2013). The intensive use of air shipping by online retail platforms for a very wide range of products shown in our paper suggests that the choice of mode of transport can also be determined by a strategy to counteract the immediacy advantage of local offline and online stores. Finally, our paper is also related to the lumpiness of trade literature (Alessandria et al.,, 2010; Kropf and Sauré,, 2014; Hornok and Koren, 2015a, ; Hornok and Koren, 2015b, ; Békés et al.,, 2017). These papers argue that there is a trade-off between per-shipment trade costs and shipping frequency. The staggering increase in the volume of low-value transactions uncovered by this paper suggests that some firms have been able to minimize the per-shipment trade costs.
@@ 1 Data and stylized facts
We obtain trade data on goods for the period 1997-2023 from the Customs and Excise Department of the Spanish Tax Agency (AEAT-Customs). Each observation captures the exchange of a product between a Spanish trade operator (a firm or an individual) and a foreign trade operator. It provides information on the direction of the trade flow (exports or imports), the Spanish trade operator’s customs identification code, the product in the 8-digit Combined Nomenclature (CN) classification, the value (in euros), the country of destination or origin, the Spanish province of destination or origin, the month in which the transaction occurred, and the date the transaction was documented in customs. It does not provide an identification of the foreign partner.
The Spanish trade data are suitable for analyzing the evolution of low-value transactions because, in contrast to other countries’ databases, it does not apply a minimum value threshold to include a transaction into the database. We remove all transactions that do not involve the transfer of property in exchange for financial compensation, such as gifts, donations, or return of goods.
@@
The AEAT-Customs database combines information from two sources: Intrastat and Extrastat. Intrastat captures firms’ trade flows with EU partners. In 2023, firms had to present a monthly Intrastat declaration if they accumulated an export (import) value with EU members of 400,000 euros in the current year or in the previous year. Extrastat captures firms’ trade flows with non-EU countries. This information is based on the customs declaration EU-based firms must present every time they perform an export or import operation with a non-EU country. Firms must present a customs declaration regardless of the amount of trade they accumulated in the previous year or have accumulated in the current year. Due to Intrastat’s data presentation threshold, AEAT-Customs does not capture all the small- and large-value transactions occurring between Spanish firms and their EU partners. This should not affect our estimations, because the below-the-Intrastat threshold trade only represented 1.8% of total trade. [^5] This percentage, estimated by AEAT-Customs, corresponds to 2022 (the last year with definitive data). It is available at: https://sede.agenciatributaria.gob.es/Sede/estadisticas/estadisticas-comercio-exterior/principales-resultados-serie/2023.html. In any case, the Online Appendix shows that there are no major differences in the evolution of the number of low-value trade transactions between Spain and EU countries and Spain and non-EU countries.
@@ C. Top 20 HS 2-digit chapters by share in total number of low-value transactions (2023)
The first stylized fact, whose documentation constitutes the main contribution of this paper, is the large increase in the share of low-value transactions in the total number of international trade transactions. It rises from 9% to 61% in exports and from 14% to 54% in imports. That is, in 2023, low-value transactions represented more than half of the total number of trade transactions. [^6] The median values of an export and import transaction in 1997 were 4,678 and 2,776 euros, respectively. In 2023, they decreased to 77 and 114 euros. As shown in Figure A.1 in the Online Appendix, this stylized fact is robust to using a low-value threshold adjusted for inflation, a low-value threshold based on the 800 USD de minimis value applied by the United States, and to splitting the sample into EU and non-EU members. [^7] The exception is that the share of low-value export transactions in the total number of transactions between 2021 and 2023 decreased in non-EU27 countries and increased in EU27 countries.
@@
The total value of low-value exports increased from 16 million euros in 1997 to 1,478 million euros in 2023 (multiplied by 92). For imports, the value of low-value transactions increased from 30 million euros to 735 million euros (multiplied by 24). These increases were much larger than those in total value, which multiplied by 4 for both exports and imports between 1997 and 2023. Despite these large relative increases, our second stylized fact is that low-value transactions still represent a small share of the total value of trade: 0.4% and 0.2% of the total value of exports and imports in 2023, respectively (panel B). [^8] In 1997, the figures were 0.02% and 0.03% for exports and imports, respectively. Fajgelbaum and Khandelwal, (2024) calculated that de minimis imports represented 7.2% of US imports of consumer goods in 2023. If we used the 800 USD threshold to define low-value transactions, these would represent 2.6% of the Spanish imports of consumer goods in 2023. For exports, the share of low-value transactions would increase to 3.8%. Therefore, when using the US de minimis threshold, low-value transactions represent a nonnegligible share of the Spanish trade in consumer goods.
@@
A major limitation of our dataset is that most low-value transactions do not report the traded product. Specifically, 70% and 71% of low-value trade transactions in 2023 did not report the exported or imported product, respectively. This occurs because export transactions equal to or below the de minimis value of the destination country, and import transactions equal to or below 150 euros are not obliged to report the traded product. Keeping this limitation in mind, the third stylized fact is that clothing and footwear represented the largest share of low-value transactions that recorded the traded product in 2023: 41% and 30% of the total number of low-value export and import transactions, respectively (panel C of Figure 1). By broad economic activities, consumption goods represented 67% and 53% of the total number of low-value export and import transactions, respectively, in 2023; intermediate goods represented 26% and 37% of export and import transactions, respectively; and parts related to capital goods represented 7% and 10% of export and import transactions, respectively. Interestingly, in 1997, books were the top product in exports and electrical machinery in imports (panel A of Figure A.2 in the Online Appendix). The share of clothing and footwear in 1997 was 12% and 13% for exports and imports, respectively. This comparison highlights that the clothing and footwear industries played an important role in the growth of the number of low-value transactions. [^9] Panel A of Figure A.3 in the Online Appendix shows that clothing and footwear also occupied prominent roles among the most exported and imported products in high-value transactions in 2023. However, their share in the total number of transactions was less than that in low-value transactions: 23% in large-value export transactions versus 41% in low-value transactions, and 22% versus 30% in import transactions.